1. Store from previous shift ended at 1.45. One hour later RHIC delivered a new nice store which ended in a MAJOR POWER DIP, fire alarm, a lot of mess in MCR. 2. Data runs: run 12728 - FS@12deg@1219B - 21k tr.2 MRS@45deg@1050B - 365k tr3. run 12731 - FS@12deg@1219B - 64k tr.2 MRS@45deg@1050B - 307k tr3. Pedestals: 12729 and 12730. 12731 ended rather abruptly with DAQ in the bloody red color. 3. Power dip caused FFS1440 and BFS1440 to trip. FFS was easy to recover, BFS got back after 1 hour. Magnets are in much worse condition now and the Magnet Crew is too busy to come over and fix it. I was joking about the floppy disk, it's on the desk. !!! No beam expected before noon !!! _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Tue Feb 1 08:00:14 2005
This archive was generated by hypermail 2.1.8 : Tue Feb 01 2005 - 08:00:37 EST